OpenStudy (anonymous):

Can someone help me with part C? A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A: Write an expression to show the total number of candies distributed by the company in x schools. (4 points) Part B: What would x(x + 1) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (anonymous):

So we have X schools. Each school has X+1 classes. So we have \[x(x+1)=x^{2}+x\] classes. Each class has 3 more students than the number of classes so: \[x^{2}+x+3\] students in each class for a total of \[(x^{2}+x)classes \times (x^{2}+x+3) \frac{students}{class}\] students. This gives us:\[x^{4}+x^{3}+3x^{2}+x^{3}+x^{2}+3x=x^{4}+2x^{3}+4x^{2}+3x\] students.